In the following situation we want to compare the mean responses in several populations.

Are you a morning person, an evening person, or neither? Does this personality trait affect how well you perform? A sample of 93 students took a psychological test that found 11 morning people, 28 evening people, and 54 who were neither. All the students then took a test of their ability to memorize at 8 a.m. and again at 9 p.m. You analyze the score at 8 a.m. minus the score at 9 p.m.

Identify the population. (choose 1)

students, faculty, staff

students that take the test   

morning people and evening people morning people, evening people

people who are neither


Identify the response variable. (choose 1)

difference in memorization score

time of day   

morning memorization score

evening memorization score


Give k, the ni, and N.

k =
n1 =
n2 =
n3 =
N =


Give the degrees of freedom of the ANOVA F test.

numerator df =
denominator     df =
Answer #1

The population:

students, faculty, staff

The response variable:

difference in memorization score

----------------

Since we divide the people in three different groups so k=3

Morning people group have 11 people so n1=11

Evening people group have 28 people so n2=28

Neither group have 54 people so n3=54

Total number of people is 93 so N=93

-------------------------

numerator df=k-1=3-1=2

denominator df=N-k=93-3=90